Saharra has 7 keys, 2 for the house and 1 for a locker. Whats the probability of not picking house or locker keys?


1
Expert's answer
2021-11-09T11:20:48-0500

"P=\frac{7-2-1}{7}=\frac{4}{7}\approx 0.5714."